The sum of two numbers is 80. The second is 8 more than 3 times the first. What are the two numbers?
Answer by checkley75(3666)

x+(3x+8)=80
4x=80-8
x=72/4
x=18 answer for the smaller number.
3*18+8=54+8=62 the larger number.
proof
18+62=80
80=80